how do you find the length of a side when you are only given a 48inch perimeter, and the length is twice the width?
Answer by rapaljer(4671)

Let x= width of the rectangle
2x= length of the rectangle

2W + 2L = P
2(x) + 2(2x)=48
2x+4x=48
6x=48
x= 8 inches Width
2x = 16 inches Length

Check:
2*8+2*16
16+ 32
48 Checks!!

For additional explanation on Perimeter problems and other Word Problems, please see my own website. Do a "Bing" search for my last name "Rapalje". Look for "Rapalje Homepage" near the top of the search list. On my Homepage near the top, look for the link "Basic, Intermediate and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in Chapter 1 for Section 1.10 Part I. In this section and others like it, you will find my own non-traditional explanations, that my own students all said were a LOT easier to understand than traditional textbooks. See also the corresponding "MATH IN LIVING COLOR" pages in which problems are solved IN COLOR!

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus